# On the Power Saving in High-Speed Ethernet-based Networks for Supercomputers and Data Centers

### v1

@@ 2.6 BXIv3 network technology
BXIv3 [15] is the network technology being developed by Eviden for HPC systems and data centers. The main objective for Eviden is to develop an interconnect that can power post-exascale systems. The existing versions of BXI, namely V1.2 [12] and V2 [6], are already deployed in several TOP500 supercomputers, including CEA-HF and Tera-1000-2 [35]. Choosing Ethernet as a basis means that there is a grounded work on standards that can speed up the development of the technology. Some examples of this set of standards are the 802.1Qbb or priority flow control (PFC) [39] for lossless per-queue link-level flow control, or the possibility to use variable-sized packets and jumbo frames.
BXIv3 network interfaces (NICs) support a link bandwidth of $400$ Gbps (expected to be $800$ Gbps in the near future), with a target deployment scale of up to 100,000 nodes, following the modular supercomputing architecture (MSA) specification [33]. BXIv3 supports different network topologies, such as Fat-trees and Megafly/Dragonfly+ [16, 31], which permit the desired levels of interconnection scalability. BXIv3 switches assume a Combined Input-Output Queued (CIOQ) switch architecture, similar to other Ethernet-based options. Regarding power consumption, we are not aware of any proposal tailored to BXIv3222Further details on the BXIv3 architecture can be obtained from one of the papers describing the BXIv3 architecture in the RED-SEA project [14].. We believe that, given the high consideration for power consumption, it will be a cornerstone for future interconnection technologies.
@@ 3.1 Link power-levels and transition logic
Fig. 5 shows the request-response mechanism that two ports from different devices, connected by a link, use to change states at the same time. At $T_{0}$, port A requests synchronization to either power up or down. Port B receives the request at $T_{1}$, and responds at $T_{2}$ with the answer. Port A has not only the response, but can also calculate the Round-Trip-Time (RTT) to establish the start of the transition, marked in the figure at ‘Sync’. Port A sends a message to signal the end of synchronization so that Port B can also calculate the RTT.
@@ 4 Evaluation
In this section, we will evaluate the mechanisms previously proposed, which we have implemented in SAURON [37], our OmNet-based network simulator. Using this simulator, we have prepared some experiments to test power-saving strategies. Regarding the network traffic, our simulator employs the VEF-TraceLib [2, 3, 5] library to inject traces from real applications into the simulator. TraceLib also provides us with a statistic of CPU usage, which will help us model the dynamic CPU usage-to-power conversion. We assume that the power consumption for nodes is mostly determined by the CPU, whose power is a direct function of usage, with a minimum power draw at idle. We also use the TopGen [36] library and its file generator to model large-scale networks with multiple topologies.
The scenario we have chosen for the experiments is a Megafly [16] topology composed of 4160 computing nodes. The network is arranged in 65 groups of 64 nodes, each group employing 16 switches with a radix of 16 (1040 switches in total). The inter-group connectivity is as described in the original Megafly publication, with one global link containing every pair of groups. Our decision on this topology is its scalability, flexibility, and popularity in HPC environments [23]. This topology can exploit the relative locality of nodes located in the same group for fast efficient communication in disjoint network partitions that exchange their messages through high-bandwidth global channels. This node arrangement can help perform partial calculations in the same-group nodes and use the global links for reductions. AI workloads, and most recently, Large Language Models (LLMs), can also benefit from this topology, as several kinds of parallelization employ different broadcast domains. Compared to the Dragonfly, the main advantages are the use of low-radix switches and the inherent absence of deadlocks. The routing algorithm employed is deterministic and minimal. When a packet has its source and destination in the same group, the algorithm behaves like D-mod-k [38]. However, when the destination is in another group, the packet travels to the spine switch that will lead to the destination group through a global link. When the packet arrives at the destination group, the downward path is the same as in D-mod-k, with only one possible route.
@@
In Table 6, we define the power parameters for each power-saving state, such as the power reduction and the time to transition from the power-saving state to the active state ($t_{w}$), and its counterpart ($t_{s}$), derived from EEE (which we showed in Tab. 3). The table also shows the percentage of power that the system devotes to the network upon applying those values to our selected scenario. The calculi for the power contribution of different elements is similar to the one by [4].
@@
Table 6: Network contribution to total power consumption in the system across two scenarios.
For the evaluation, we have selected traces of two HPC applications (LAMMPS, PATMOS) and two distributed Deep Learning training sessions (MLWF [22] and AlexNet [21]) to study the network behavior in both contexts. For each application, we will first describe the behavior of the application and the potential for energy saving. First, we will compare how much energy we save with different techniques in comparison to not using any power-saving technique. Non-energy metrics include the degradation that is introduced by power-saving strategies. Mainly, we want to focus on the increase in execution time and latency. Specifically, the latency metric we are exploring is the packet latency, which measures the time from the packet generation to its arrival at the destination.
@@ 4.1 Evaluation for the LAMMPS application
The LAMMPS molecular dynamics uses the network to exploit its parallel architecture and k.pdf the model distributed among nodes. The simulation space is divided into subdomains assigned to different processors, each of which manages own atoms and ghost atoms—copies of neighboring atoms needed to calculate interactions.
@@
In summary, LAMMPS creates a mix of localized and global communication patterns, with MPI_AllReduce identified as the primary contributor to potential network bottlenecks, especially in large-scale simulations. Being an all-to-all communication operation, MPI_AllReduce will use most paths between nodes. This means that this operation will have the most degradation if all links are in a power-saving state. Fig. 6 shows the network efficiency or throughput (normalized against the total network bandwidth) for the complete execution time (2.431 seconds) in the configured system. The total energy consumed is 9.993 MJ by nodes and 1.846 MJ by the network, totaling 11.839 MJ for the whole system. We can see that after a second of computation, the network activity starts being intermittent with several traffic spikes due to the inter-node collective operations.
@@ 4.1.1 Results with fixed $t_{PDT}$ values
The first statistic we will examine is the execution time, displayed in Fig. 7(a). We can see that using Deep Sleep causes the execution time to go beyond 100 %, while Fast Wake never surpasses 10 % overhead with $t_{PDT}$ values below 10 µs. This coincides with the $t_{w}$ for Deep Sleep being an order of magnitude above that of Fast Wake. We can see the overhead drops significantly if a $t_{PDT}$ of 100 µs or higher is chosen, as this interval surpasses the length for most of the inactivity periods.
@@
Fig. 7 shows the impact of different $t_{PDT}$ values on execution time, energy savings, and packet latency for LAMMPS. As we can see in Fig. 7(b), Deep Sleep achieves negative power consumption reduction when using 10 µs and below, in accordance with the increment in execution time shown in Fig. 7(a). We can also see that using Fast Wake produces an increment of power consumption with $t_{PDT}$ below 1 µs, but it achieves energy savings close to 10% with $t_{PDT}$ values larger than 10 µs. In this latter case, Fast Wake obtains an overhead in the execution time of less than 10%. With more permissive $t_{PDT}$ values, Deep Sleep produces more power savings.
If we look at the packet latency overhead in Fig. 7(c), we can see that the tendency for each $t_{PDT}$ value is similar to that of the execution time, albeit on another order of magnitude.
@@ 4.1.2 Results using PerfBound and PerfBoundCorrect
As shown in Fig. 8(a), we can see the overhead on the application when using PerfBound and PerfBoundCorrect techniques. Just like with a fixed $t_{PDT}$, the difference between Fast Wake and Deep Sleep remains in the same proportion. It is important to note that using either of the histogram types makes a negligible difference in the case of Fast Wake, but has an effect when using Deep Sleep. Using a circular buffer to store the inactivity periods provides the worst results by a noticeable margin. We can also see that clearing the histogram instead of replacing old records is better for the execution time, likely because values become more permissive each time the histogram is reset. Recording all the inactivity periods provides a lower execution time overhead. However, the advantage of the self-clearing histogram comes at the cost of storing an indefinite number of bins, which is not feasible.
@@
The overhead on execution time has a direct impact on the energy consumed, as seen in Fig. 8(b). Both PerfBound and PerfBoundCorrect provide similar levels of energy savings, with slightly higher values using PerfBoundCorrect. While PerfBound with Deep Sleep does not lead to energy savings, for thresholds of 1 % and 2 % PerfBoundCorrect manages to revert that situation. When using a high threshold, however, energy saving does not occur because the calculated $t_{PDT}$ values are too short.
Finally, Fig. 8(c) shows that PerfBoundCorrect, in all cases and with either power-saving state, reduces latency significantly. It is clear that, particularly when using Deep Sleep, a circular buffer to control the size of the histogram produces the highest overhead.
@@ 4.2 Evaluation for the PATMOS application
PATMOS is a Monte Carlo neutron transport code developed at the CEA (France), designed to prototype applications in nuclear safety and radiation shielding [9].
@@
This last description coincides with the evolution of network efficiency shown in Fig. 9. As we can see, the application only accesses the network during startup and upon completion. The rest of the time, as nodes are executing independent simulations, is an opportunity for power saving on the network. With no power-saving strategy on links, the application execution time is 1290.64 seconds. During that time, the whole system consumed 7.334 GJ, operating at 5.685 MW, very close to maximum cluster capacity. The network has consumed 345.28 KJ. In terms of efficiency, the whole system is at 8,75E mJ/bit, and the network is at 412,08 nJ/bit.
@@ 4.2.1 Results with fixed ${t}_{PDT}$ values
The other possible overhead for the trace is related to latency. If we look at Fig. 10(a), we can see the effect each $t_{PDT}$ value has on packet latency. As we can see, employing Deep Sleep increases the packet latency by up to 7.5 %, with a steep increase on the microsecond scale and below.
@@
Regarding the total energy consumed, leaving links operational after transmission is potentially beneficial for future outgoing packets at the cost of more energy consumed over time. Fig. 10(b) shows the effect each PDT value has on the total energy consumed by the system. As is the case with execution time, the ports experience few transitions because of the traffic pattern of the application. As such, the energy saved on links is very high, with ports being, on average, on for only 0.02 % of the time.
@@ 4.2.2 Results using PerfBound and PerfBoundCorrect
If we look at Fig. 11(b), we can see the total energy savings on the system due to applying the power-saving techniques on the network links. We can see that for this application, there is no relevant difference in using one histogram type over the other, but using Deep Sleep yields greater energy savings. As expected in a situation with links having few and long inactivity periods, Deep Sleep provides better power saving than Fast Wake.
@@ 4.3 Evaluation for the MLWF application
The MLWF (Machine Learning Weather Forecast) [22] application is developed within the context of the MAELSTROM European project. The trace was collected during a training session using Horovod [29]. The traffic pattern in this application, in particular, and in Deep Learning training sessions in general, usually involves the usage of MPI collective operations for data exchange. Particularly, the most predominant traffic portion comes from AllReduce operations. Usually, every layer of the neural network that is being trained runs a repetitive sequence of Gather-Gatherv-Broadcast-Broadcast until the weights converge, then the AllReduce is invoked to have the information on every node [5].
The network efficiency results for this trace are shown in Fig. 12. We have opted to show them on a logarithmic scale to illustrate that any traffic rate, regardless of the value, provides great difficulty in finding opportunities for power saving.
@@ 4.3.1 Results with fixed ${t}_{PDT}$ values
Fig. 13(a) shows the increase in execution time for the application on different experiments. We can see that for small $t_{PST}$ values, the overhead is in consonance with the $t_{w}$ values for Deep Sleep and Fast Wake. To halve worst-case execution time overhead using Deep Sleep, we would have to set a $t_{PDT}$ value above 1 us, and another 3 orders of magnitude to reduce it to a third.
@@
As shown in Fig. 13(b), we can see that there is an evident difference between the latency overhead that occurs with either power-saving state. As we increase $t_{PDT}$ values, the latency starts dropping to lower, more acceptable values.
The results for energy saving for this application are shown in Fig. 13(c). As shown, $t_{PDT}$ values of 10 and 100 µs increase the energy saved in both power-saving states. Lower values provide marginal power-saving values or even increased energy usage in that proportion. Indeed, even if the traffic between the large bursts at the start and end of the applications, there is communication that is not negligible for us. However, once we use higher values, the energy reduction begins to diminish. This is due to inactivity, 99 % of periods being within the millisecond range. This is also the reason why using a 1-second $t_{PDT}$ provides more power saving than a millisecond-scale one.
@@ 4.3.2 Results using PerfBound and PerfBoundCorrect
Fig. 14(a) shows the overhead on execution time when using PerfBound and PerfBoundCorrect with either power-saving state on links. We can see that using Fast Wake impacts the execution time by less than 1 %, while the overhead is higher on Deep Sleep, due to its $t_{w}$ value. Still, we can see that PerfBoundCorrect reduces the overhead compared to PerfBound.
Regarding energy savings results displayed in Fig. 14(b), we can see that for this application, PerfBoundCorrect provides around 2 % less energy savings than PerfBound when employing Deep Sleep, and a smaller difference on the Fast Wake experiments. The results for energy consumed show similar trends as those for execution time (Fig. 14(a)), meaning that for this application, having more permissive $t_{PDT}$ values simply increments the time that links are on. This is because PerfBound already hits most of its predictions.
The packet latency results are shown in Fig. 14(c). We can see that for both Deep Sleep and Fast Wake, the overhead is very small. This further proves that PerfBound already makes good predictions, and PerfBoundCorrect produces diminishing returns.

### v2

@@ 2.6 BXIv3 network technology
BXIv3 [bxi3eviden] is the network technology being developed by Eviden for HPC systems and data centers. The main objective for Eviden is to develop an interconnect that can power post-exascale systems. The existing versions of BXI, namely V1.2 [bxihoti] and V2 [bxi2atos], are already deployed in several TOP500 supercomputers, including CEA-HF and Tera-1000-2 [top500]. Choosing Ethernet as a basis means that there is a grounded work on standards that can speed up the development of the technology. Some examples of this set of standards are the 802.1Qbb or priority flow control (PFC) [cclsrdmad] for lossless per-queue link-level flow control, or the possibility to use variable-sized packets and jumbo frames.
BXIv3 network interfaces (NICs) support a link bandwidth of $400$ Gbps (expected to be $800$ Gbps in the near future), with a target deployment scale of up to 100,000 nodes, following the modular supercomputing architecture (MSA) specification [Suarez19MSA]. BXIv3 supports different network topologies, such as Fat-trees and Megafly/Dragonfly+ [flajslik_megafly_2018, dfly+], which permit the desired levels of interconnection scalability. BXIv3 switches assume a Combined Input-Output Queued (CIOQ) switch architecture, similar to other Ethernet-based options. Regarding power consumption, we are not aware of any proposal tailored to BXIv3222Further details on the BXIv3 architecture can be obtained from one of the papers describing the BXIv3 architecture in the RED-SEA project [Gomez24redsea].. We believe that, given the high consideration for power consumption, it will be a cornerstone for future interconnection technologies.
@@ 3.1 Link power-levels and transition logic
Figure 5 shows the request-response mechanism that two ports from different devices, connected by a link, use to change states at the same time. At $T_{0}$, port A requests synchronization to either power up or down. Port B receives the request at $T_{1}$, and responds at $T_{2}$ with the answer. Port A has not only the response, but can also calculate the Round-Trip-Time (RTT) to establish the start of the transition, marked in the figure at ‘Sync’. Port A sends a message to signal the end of synchronization so that Port B can also calculate the RTT.
@@ 4 Evaluation
In this section, we will evaluate the mechanisms previously proposed, which we have implemented in SAURON [sauron], our OmNet-based network simulator. Using this simulator, we have prepared some experiments to test power-saving strategies. Regarding the network traffic, our simulator employs the VEF-TraceLib [VEF-traces, Andujar16JSC, andujar_extending_2023] library to inject traces from real applications into the simulator. TraceLib also provides us with a statistic of CPU usage, which will help us model the dynamic CPU usage-to-power conversion. We assume that the power consumption for nodes is mostly determined by the CPU, whose power is a direct function of usage, with a minimum power draw at idle. We also use the TopGen [topgen] library and its file generator to model large-scale networks with multiple topologies.
The scenario we have chosen for the experiments is a Megafly [flajslik_megafly_2018] topology composed of 4160 computing nodes. The network is arranged in 65 groups of 64 nodes, each group employing 16 switches with a radix of 16 (1040 switches in total). The inter-group connectivity is as described in the original Megafly publication, with one global link containing every pair of groups. Our decision on this topology is its scalability, flexibility, and popularity in HPC environments [mflyadapt]. This topology can exploit the relative locality of nodes located in the same group for fast efficient communication in disjoint network partitions that exchange their messages through high-bandwidth global channels. This node arrangement can help perform partial calculations in the same-group nodes and use the global links for reductions. AI workloads, and most recently, Large Language Models (LLMs), can also benefit from this topology, as several kinds of parallelization employ different broadcast domains. Compared to the Dragonfly, the main advantages are the use of low-radix switches and the inherent absence of deadlocks. The routing algorithm employed is deterministic and minimal. When a packet has its source and destination in the same group, the algorithm behaves like D-mod-k [PGFT]. However, when the destination is in another group, the packet travels to the spine switch that will lead to the destination group through a global link. When the packet arrives at the destination group, the downward path is the same as in D-mod-k, with only one possible route.
@@
In Table 6, we define the power parameters for each power-saving state, such as the power reduction and the time to transition from the power-saving state to the active state ($t_{w}$), and its counterpart ($t_{s}$), derived from EEE (which we showed in Table 3). The table also shows the percentage of power that the system devotes to the network upon applying those values to our selected scenario. The calculi for the power contribution of different elements is similar to the one by [ANDUJAR2023126].
@@
Table 6: Network contribution to total system power consumption across two scenarios.
For the evaluation, we have selected traces of two HPC applications (LAMMPS, PATMOS) and two distributed Deep Learning training sessions (MLWF [MLWF] and AlexNet [ALEXNET]) to study the network behavior in both contexts. For each application, we will first describe the behavior of the application and the potential for energy saving. First, we will compare how much energy we save with different techniques in comparison to not using any power-saving technique. Non-energy metrics include the degradation that is introduced by power-saving strategies. Mainly, we want to focus on the increase in execution time and latency. Specifically, the latency metric we are exploring is the packet latency, which measures the time from the packet generation to its arrival at the destination.
@@ 4.1 Evaluation for the LAMMPS application
The LAMMPS molecular dynamics uses the network to exploit its parallel architecture and keeps the model distributed among nodes. The simulation space is divided into subdomains assigned to different processors, each of which manages own atoms and ghost atoms—copies of neighboring atoms needed to calculate interactions.
@@
In summary, LAMMPS creates a mix of localized and global communication patterns, with MPI_AllReduce identified as the primary contributor to potential network bottlenecks, especially in large-scale simulations. Being an all-to-all communication operation, MPI_AllReduce will use most paths between nodes. This means that this operation will have the most degradation if all links are in a power-saving state. Figure 6 shows the network efficiency or throughput (normalized against the total network bandwidth) for the complete execution time (2.431 seconds) in the configured system. The total energy consumed is 9.993 MJ by nodes and 1.846 MJ by the network, totaling 11.839 MJ for the whole system. We can see that after a second of computation, the network activity starts being intermittent with several traffic spikes due to the inter-node collective operations.
@@ 4.1.1 Results with fixed $t_{PDT}$ values
The first statistic we will examine is the execution time, displayed in Figure 7(a). We can see that using Deep Sleep causes the execution time to go beyond 100 %, while Fast Wake never surpasses 10 % overhead with $t_{PDT}$ values below 10 µs. This coincides with the $t_{w}$ for Deep Sleep being an order of magnitude above that of Fast Wake. We can see the overhead drops significantly if a $t_{PDT}$ of 100 µs or higher is chosen, as this interval surpasses the length for most of the inactivity periods.
@@
Figure 7 shows the impact of different $t_{PDT}$ values on execution time, energy savings, and packet latency for LAMMPS. As we can see in Figure 7(b), Deep Sleep achieves negative power consumption reduction when using 10 µs and below, in accordance with the increment in execution time shown in Figure 7(a). We can also see that using Fast Wake produces an increment of power consumption with $t_{PDT}$ below 1 µs, but it achieves energy savings close to 10% with $t_{PDT}$ values larger than 10 µs. In this latter case, Fast Wake obtains an overhead in the execution time of less than 10%. With more permissive $t_{PDT}$ values, Deep Sleep produces more power savings.
If we look at the packet latency overhead in Figure 7(c), we can see that the tendency for each $t_{PDT}$ value is similar to that of the execution time, albeit on another order of magnitude.
@@ 4.1.2 Results using PerfBound and PerfBoundCorrect
As shown in Figure 8(a), we can see the overhead on the application when using PerfBound and PerfBoundCorrect techniques. Just like with a fixed $t_{PDT}$, the difference between Fast Wake and Deep Sleep remains in the same proportion. It is important to note that using either of the histogram types makes a negligible difference in the case of Fast Wake, but has an effect when using Deep Sleep. Using a circular buffer to store the inactivity periods provides the worst results by a noticeable margin. We can also see that clearing the histogram instead of replacing old records is better for the execution time, likely because values become more permissive each time the histogram is reset. Recording all the inactivity periods provides a lower execution time overhead. However, the advantage of the self-clearing histogram comes at the cost of storing an indefinite number of bins, which is not feasible.
@@
The overhead on execution time has a direct impact on the energy consumed, as seen in Figure 8(b). Both PerfBound and PerfBoundCorrect provide similar levels of energy savings, with slightly higher values using PerfBoundCorrect. While PerfBound with Deep Sleep does not lead to energy savings, for thresholds of 1 % and 2 % PerfBoundCorrect manages to revert that situation. When using a high threshold, however, energy saving does not occur because the calculated $t_{PDT}$ values are too short.
Finally, Figure 8(c) shows that PerfBoundCorrect, in all cases and with either power-saving state, reduces latency significantly. It is clear that, particularly when using Deep Sleep, a circular buffer to control the size of the histogram produces the highest overhead.
@@ 4.2 Evaluation for the PATMOS application
PATMOS [PATMOS] is a Monte Carlo neutron transport code developed at the CEA (France), designed to prototype applications in nuclear safety and radiation shielding.
@@
This last description coincides with the evolution of network efficiency shown in Figure 9. As we can see, the application only accesses the network during startup and upon completion. The rest of the time, as nodes are executing independent simulations, is an opportunity for power saving on the network. With no power-saving strategy on links, the application execution time is 1290.64 seconds. During that time, the whole system consumed 7.334 GJ, operating at 5.685 MW, very close to maximum cluster capacity. The network has consumed 345.28 KJ. In terms of efficiency, the whole system is at 8,75E mJ/bit, and the network is at 412,08 nJ/bit.
@@ 4.2.1 Results with fixed ${t}_{PDT}$ values
The other possible overhead for the trace is related to latency. If we look at Figure 10(a), we can see the effect each $t_{PDT}$ value has on packet latency. As we can see, employing Deep Sleep increases the packet latency by up to 7.5 %, with a steep increase on the microsecond scale and below.
@@
Regarding the total energy consumed, leaving links operational after transmission is potentially beneficial for future outgoing packets at the cost of more energy consumed over time. Figure 10(b) shows the effect each PDT value has on the total energy consumed by the system. As is the case with execution time, the ports experience few transitions because of the traffic pattern of the application. As such, the energy saved on links is very high, with ports being, on average, on for only 0.02 % of the time.
@@ 4.2.2 Results using PerfBound and PerfBoundCorrect
If we look at Figure 11(b), we can see the total energy savings on the system due to applying the power-saving techniques on the network links. We can see that for this application, there is no relevant difference in using one histogram type over the other, but using Deep Sleep yields greater energy savings. As expected in a situation with links having few and long inactivity periods, Deep Sleep provides better power saving than Fast Wake.
@@ 4.3 Evaluation for the MLWF application
The MLWF (Machine Learning Weather Forecast) [MLWF] application is developed within the context of the MAELSTROM European project. The trace was collected during a training session using Horovod [sergeev2018horovod2]. The traffic pattern in this application, in particular, and in Deep Learning training sessions in general, usually involves the usage of MPI collective operations for data exchange. Particularly, the most predominant traffic portion comes from AllReduce operations. Usually, every layer of the neural network that is being trained runs a repetitive sequence of Gather-Gatherv-Broadcast-Broadcast until the weights converge, then the AllReduce is invoked to have the information on every node [andujar_extending_2023].
The network efficiency results for this trace are shown in Figure 12. We have opted to show them on a logarithmic scale to illustrate that any traffic rate, regardless of the value, provides great difficulty in finding opportunities for power saving.
@@ 4.3.1 Results with fixed ${t}_{PDT}$ values
Figure 13(a) shows the increase in execution time for the application on different experiments. We can see that for small $t_{PST}$ values, the overhead is in consonance with the $t_{w}$ values for Deep Sleep and Fast Wake. To halve worst-case execution time overhead using Deep Sleep, we would have to set a $t_{PDT}$ value above 1 us, and another 3 orders of magnitude to reduce it to a third.
@@
As shown in Figure 13(b), we can see that there is an evident difference between the latency overhead that occurs with either power-saving state. As we increase $t_{PDT}$ values, the latency starts dropping to lower, more acceptable values.
The results for energy saving for this application are shown in Figure 13(c). As shown, $t_{PDT}$ values of 10 and 100 µs increase the energy saved in both power-saving states. Lower values provide marginal power-saving values or even increased energy usage in that proportion. Indeed, even if the traffic between the large bursts at the start and end of the applications, there is communication that is not negligible for us. However, once we use higher values, the energy reduction begins to diminish. This is due to inactivity, 99 % of periods being within the millisecond range. This is also the reason why using a 1-second $t_{PDT}$ provides more power saving than a millisecond-scale one.
@@ 4.3.2 Results using PerfBound and PerfBoundCorrect
Figure 14(a) shows the overhead on execution time when using PerfBound and PerfBoundCorrect with either power-saving state on links. We can see that using Fast Wake impacts the execution time by less than 1 %, while the overhead is higher on Deep Sleep, due to its $t_{w}$ value. Still, we can see that PerfBoundCorrect reduces the overhead compared to PerfBound.
Regarding energy savings results displayed in Figure 14(b), we can see that for this application, PerfBoundCorrect provides around 2 % less energy savings than PerfBound when employing Deep Sleep, and a smaller difference on the Fast Wake experiments. The results for energy consumed show similar trends as those for execution time (Figure 14(a)), meaning that for this application, having more permissive $t_{PDT}$ values simply increments the time that links are on. This is because PerfBound already hits most of its predictions.
The packet latency results are shown in Figure 14(c). We can see that for both Deep Sleep and Fast Wake, the overhead is very small. This further proves that PerfBound already makes good predictions, and PerfBoundCorrect produces diminishing returns.